# SURGE: On the Potential of Large Language Models as General-Purpose Surrogate Code Executors

### v3

@@ Abstract
Neural surrogate models have emerged as powerful and efficient tools in data mining. Meanwhile, large language models (LLMs) have demonstrated remarkable capabilities in code-related tasks. We investigate a novel application: using LLMs as surrogate models for code execution prediction. Given LLMs’ unique ability to understand and process diverse programs, they present a promising direction for building general-purpose surrogate models. To systematically investigate this capability, we introduce SURGE, a comprehensive benchmark with problems covering key aspects: multi-language programming tasks, competition-level programming problems, repository-level code analysis, high-cost scientific computing, time-complexity-intensive algorithms, buggy code analysis, programs dependent on specific compilers or execution environments, and formal mathematical proof verification. Through extensive empirical analysis of open-source and proprietary LLMs, we examine scaling laws, data efficiency, and predictive accuracy. Our findings reveal important insights about the feasibility of LLMs as efficient surrogates for computational processes, with implications for automated software testing, program analysis, and computational resource optimization in data mining applications. Code and dataset are released at https://github.com/Imbernoulli/SURGE.
@@ 1 Introduction
Neural surrogate models (Zhang et al., 2024; Sun and Wang, 2019) have emerged as powerful tools in data mining and machine learning, efficiently approximating complex computational processes. Meanwhile, Large language models (LLMs) (Reid et al., 2024; Meta, 2024; Anthropic, 2024b; Hui et al., 2024; Bi et al., 2024) have demonstrated remarkable capabilities in code-related tasks (Lu et al., 2021a; Zheng et al., 2023; Luo et al., 2023; Team, 2024a; Guo et al., 2024), including code understanding (Ahmad et al., 2020; Chakraborty et al., 2020) and code generation (Li et al., 2018a; Parvez et al., 2018). However, an equally important yet underexplored question is whether LLMs can serve as general-purpose surrogate code executors, which predict the output and behavior of a program without actually running it. Recent study (Lyu et al., 2024) acknowledges its importance, however, it focuses on a case study rather than a systematic analysis.
@@
The ability to predict code execution outcomes without execution has tremendous siginficance. In scientific computing, running simulations often requires substantial computational resources and is time-consuming, making it impractical to test every possible configuration (Lu and Ricciuto, 2019; Hesthaven and Ubbiali, 2018; Benner et al., 2015). In security-sensitive environments, executing untrusted code poses inherent risks, necessitating alternative mechanisms for assessing program behavior without exposing the system to potential vulnerabilities (Nebbione and Calzarossa, 2023; Shirazi et al., 2017; Wang et al., 2024). Additionally, some code requires highly specific execution environments, which may not always be available, making surrogate execution a valuable alternative (Queiroz et al., 2023; Gu et al., 2025). Moreover, accurately predicting a model’s potential outputs or errors is crucial for improving traditional tasks such as code understanding, code generation, and even math reasoning (Li et al., 2025). Lastly, many works use LLMs as reward models (RMs) in reinforcement learning. For code tasks, accurate execution prediction is key to a reliable RM (Ouyang et al., 2022).
@@
Traditional approaches to surrogate code executing (King, 1976; Cadar and Sen, 2013) struggle to generalize across languages and suffer from scalability issues when applied to complex real-world codebases. Containerized environments (Merkel, 2014) mitigate dependency issues but still require full code execution. Recent efforts to train neural executors (Yan et al., 2020) focus on narrow tasks and lack the generality needed for real-world code. In contrast, LLMs’ capacity to internalize patterns from vast code corpora (Lu et al., 2021b; Chaudhary, 2023) suggests a path toward general-purpose surrogate code execution, though their limits remain unquantified.
Therefore, it’s crucial to understand the potential of LLMs as general-purpose surrogate code executors. In this paper, we try to figure out how well current LLMs function as general-purpose surrogate code executors and whether this ability can be improved through training.
@@
To systematically investigate LLMs as GEneral-purpose SURrogate code executors, we introduce SURGE. It includes components: (1) fundamental programming tasks in multiple languages, (2) competition programming problems requiring deep logical inference, (3) repository-level codebases that test long-range dependencies, (4) scientific simulations and optimizations where direct execution is high-cost, (5) time-consuming logical algorithms that have high time-complexity, (6) buggy code that examines LLMs’ ability to predict runtime errors, (7) programs whose behavior depends on specific compiler versions or execution environments and (8) math theorem proving in formal language (LEAN4) (De Moura et al., 2015; Moura and Ullrich, 2021) which expects compilers to testify the proofs.
@@
Through extensive empirical evaluation of 21 open-source and proprietary LLMs on SURGE, we provide the first large-scale study of LLMs’ capabilities as computational surrogates. Additionally, we investigate the impact of various factors including prompt engineering strategies, programming language characteristics, computational complexity, and execution time requirements on surrogate performance. Our findings reveal both the promising potential and current limitations of LLMs as general-purpose code execution surrogates. The performance of typical models on SURGE is shown in Figure 1.
@@
Beyond evaluating existing models, we conduct a scaling law study to determine whether models’ performance improves with increased model size and data availability. Specifically, we train models with sizes ranging from {0.5/1.5/3/7}B on different scales of training data from the formal language subset of SURGE and analyze how model capacity and training data scale affect surrogate execution accuracy. Our scaling experiments demonstrate that performance consistently improves with both model size (from 0.5B to 7B parameters) and training steps, with larger models showing stronger learning capacity and higher performance ceilings throughout the training process (Figure 2).
@@
We introduce SURGE, the first but holistic benchmark for evaluating LLMs as general-purpose surrogate code executors. It consists of subsets and problems.
We conduct an extensive evaluation of open-source and proprietary LLMs on SURGE, providing the first large-scale study assessing their capabilities.
We present a scaling law study with models of varying sizes and data of different scales, providing empirical insights into whether LLMs systematically improve with more capacity and training data.
@@ 2 Related Works
Neural surrogate models are neural network-based approximations used to replace computationally expensive simulations in various scientific and engineering domains (Zhang et al., 2024; Sun and Wang, 2019). These models act as efficient emulators by learning complex input-output relationships from high-fidelity data, significantly reducing computational costs while maintaining accuracy (Raissi et al., 2020; Sun et al., 2020; Bessa et al., 2017; Thuerey et al., 2020; Raissi et al., 2019; Willard et al., 2022). Recently, generative models (e.g. pre-trained language models) have been incorporated into surrogate modeling. Some equip language models with traditional surrogate models to facilitate iterative optimization (Ma et al., 2024; Lyu et al., 2025), and some use generative models to realize the end-to-end surrogate process (Gruver et al., 2024; Hao et al., 2024; Wimmer and Rekabsaz, 2023; Che et al., 2024). While these studies primarily focus on natural sciences, time series, and multimodal gaming, the application of surrogate modeling to code execution, where both input and output exist in the modality of language, remains unexplored.
@@
LLMs are widely used in code-related tasks (Lu et al., 2021a; Zheng et al., 2023; Luo et al., 2023; Team, 2024a; Guo et al., 2024), which can be fundamentally categorized into code understanding and code generation. Code understanding tasks include code summarization Hu et al. (2018); Harer et al. (2019); Ahmad et al. (2020), bug detection Li et al. (2018b); Russell et al. (2018); Zhou et al. (2019); Chakraborty et al. (2020), duplication detection Zhang et al. (2019); Yu et al. (2019); Wang et al. (2020), code retrieval (Husain et al., 2020; Lu et al., 2021a), etc. Code generation tasks include code completion Li et al. (2018a); Parvez et al. (2018), code repair Chen et al. (2019); Chakraborty et al. (2020); Lutellier et al. (2020), test generation (Watson et al., 2020; Siddiq et al., 2024; schäfer2023empiricalevaluationusinglarge), etc. When evaluating LLMs on these tasks, some datasets provide broad evaluations across general tasks (Chen et al., 2021; Austin et al., 2021; Liu et al., 2024; Muennighoff et al., 2023), some focus on specific dimensions such as multi-lingual capabilities (Athiwaratkun et al., 2022; Zheng et al., 2023; Cassano et al., 2022; Khan et al., 2023), competition problems (Hendrycks et al., 2021; Li et al., 2022), data science tasks (Chandel et al., 2022; Lai et al., 2023; Huang et al., 2022), repository-level understanding (Zhang et al., 2023a; Shrivastava et al., 2023; Liu et al., 2023; Li et al., 2024; Jimenez et al., 2023; Ding et al., 2024; Zan et al., 2024), and reasoning abilities (Athiwaratkun et al., 2022; Austin et al., 2021; Cobbe et al., 2021; Gao et al., 2023; Gu et al., 2024). However, while the potential execution result of code is important for both code understanding and generation, this aspect remains largely unexplored (Weber et al., 2024).
@@ 3 SURGE
To comprehensively evaluate the potential of LLMs as surrogate code executors, we construct a diverse benchmark, SURGE, that covers various code execution scenarios. Our dataset is designed to assess the model’s ability to approximate execution results across multiple dimensions, including multi-lingual diversity, repository-level complexity, computational intensity, error handling, and scenario-dependent variability. Below, we describe each component of SURGE, explaining the motivation behind its inclusion.
@@ 3.1 Dataset Construction
As illustrated in Figure 3, the construction of SURGE involves four distinct methodologies, each applied to specific components of our eight subsets. For ML, CL, BG components, we employ the Iterative Refactor methodology, where we refine the code through an interactive process involving LLM assistance and human verification. The RL component is constructed through Repository Sampling, where we extract and construct test cases from both public and custom code repositories. For SC, TC, DR components, we utilize Manual Implementation, carefully handcrafting code based on selected textbook materials and question collections. Finally, the FL component is developed using the Inference & Verification approach, leveraging formal mathematical provers to generate proofs and validate them through compiler verification. These methodologies are strategically applied to their respective components to ensure a comprehensive and rigorous benchmark construction.
@@ 3.2 Dataset Components
A fundamental nature of a general-purpose surrogate executor is its ability to handle multiple programming languages. Since we are exploring the model’s execution capabilities, we do not focus on rendering languages such as HTML, but concentrate on computational languages. Our dataset covers 7 such languages, including C, C++, C#, Java, Rust, Python, and Julia. Our dataset is adapted from McEval (Chai et al., 2024). The original dataset does not provide executable code, so we used an LLM to generate executable code by providing it with prompts, ground truth, and test cases in the original dataset. This generated code was then manually processed. We manually modified code that failed to compile, such as adding missing headers. And then we took precautions to prevent answers from being leaked in assert statements or comments.
@@
Next, we consider competition-level code, which presents a higher level of coding difficulty. We collect these tasks from 2 public repositories 111https://github.com/azl397985856/leetcode222https://gitee.com/shanire/OJCode, which contain problems from open coding platforms (e.g. LeetCode 333https://leetcode.com, Luogu 444https://www.luogu.com.cn). The dataset includes problems in 3 languages, C++, Java, and JavaScript. Since the original repositories only provide partial solutions, we first use an LLM to generate complete, executable code that prints the expected output. This generated code is then manually verified. To investigate whether problem difficulty affects the performance of surrogate models, we further employ an LLM to automatically classify problems into different difficulty levels, followed by human verification to ensure accuracy.
@@
Scientific computing has long been adopting neural surrogate models. It’s crucial to investigate whether LLMs can approximate execution results for these non-trivial tasks and hold the potential to serve as efficient surrogate models. We introduced tasks ranging from solving ordinary differential equations (ODEs) to optimization problems and signal processing. These tasks are motivated by and widely used in real-world scientific challenges, including areas where increasing research has been done on solving these computational tasks through building efficient surrogate models (Wu et al., 2023; Zhang et al., 2023b)..
@@
Surrogate models were originally motivated by real-world applications where program execution is high-cost and time-consuming. It’s a necessity for LLms to generalize well to strongly computation-power-dependent and time-consuming tasks. We include examples from linear algebra, sorting, searching, Monte Carlo simulations, and string matching programs, ensuring a broad representation of computationally intense tasks. These tasks cover various complexity classes, including P (e.g. sorting an array), NP (e.g. Hamiltonian Cycle), and NP-Hard (e.g. Traveling Salesman’s Problem). Additionally, we record the CPU execution time for each program under consistent environments individually to support subsequent studies.
@@
Real-world code execution often encounters errors due to syntax mistakes, logical flaws, or runtime exceptions. Code errors pose risks in sensitive scenarios, therefore, we aim for the code surrogate model to recognize the presence of bugs. This dataset is adapted from DebugBench (Tian et al., 2024), which extracted Java, Python, and C++ code from LeetCode and manually inserted errors from 4 major bug categories and 18 minor types. Since DebugBench only provides erroneous code snippets rather than complete executable programs, we first used an LLM to automatically complete the error-free code into fully runnable versions. After verifying their correctness, we replaced relevant parts with buggy code and executed them again to capture the corresponding error outputs. Some errors resulted in infinite loops causing timeouts, so we set a -second execution filter out such cases.
@@
Various contextual factors, such as compiler versions, optimization levels, and language standards often influence code execution. These variations can lead to different outputs for the same code snippet. It is crucial for surrogate models to recognize such discrepancies and adapt to different settings. We focus specifically on C++ and manually collect code snippets from textbooks and online sources (Bryant and O’Hallaron, 2010; Lippman et al., 2012) that exhibit different behaviors under varying compilation settings. We consider multiple compilers (g++, clang++), C++ standards (03, 11, 14, 17), and optimization levels (-O0, -O1, -O2, -O3, -Os). Each snippet is executed across these different settings, and we retain only those that produce varying outputs through different configurations while discarding cases that yield identical results across all settings.
@@
Math-Proving Formal Languages are specialized programming languages designed for mathematical proof verification through compilers (De Moura et al., 2015; Moura and Ullrich, 2021; Paulson, 1994; Barras et al., 1997). These compilers can determine whether a proof is correct and identify specific errors if present. Formal theorem proving is gaining increasing attention, as constructing a valid proof requires extensive trial and verification, which can be highly time-consuming. We aim to explore whether a surrogate model can assist in this verification process. Unlike conventional code execution, formal language verification follows a distinct paradigm, therefore we categorize this task separately. In this study, we focus on Lean 4 which is the most widely used proof assistant language. To build our dataset, we use state-of-the-art formal math prover, Goedel-Prover (Lin et al., 2025), to conduct large-scale reasoning on Lean-Workbook (Ying et al., 2024), extracting an equal proportion of correct and incorrect proofs. This balanced dataset allows us to evaluate the surrogate model’s ability to assess proof validity effectively.
| Subset | Construction Method | Source | Quantity | Example Num. | Metric | Categories |
| ML | Iterative Refactor | McEval (Chai et al., 2024) | 150 | 3 | Exact Match | Languages |
| CL | Iterative Refactor | GitHub | 150 | 3 | Exact Match | Difficulties & Languages |
| RL | Repository Sampling | GitHub & Custom | 60 | 3 | Mixed | - |
| SC | Manual Implementation | Custom | 150 | 3 | Mixed | Scenarios |
| TC | Manual Implementation | Custom | 150 | 3 | Mixed | Scenarios, CPU Time |
| BG | Iterative Refactor | DebugBench (Tian et al., 2024) | 150 | 4 | Jaccard similarity | Error Type & Languages |
| DR | Manual Implementation | Custom | 200 | 3 | Jaccard similarity | Variable Type |
| FL | Inference & Verification | Lean-Workbook (Ying et al., 2024) | 150 | 3 | Custom | - |

### v4

@@ Abstract
Neural surrogate models are powerful and efficient tools in data mining. Meanwhile, large language models (LLMs) have demonstrated remarkable capabilities in code-related tasks, such as generation and understanding. However, an equally important yet underexplored question is whether LLMs can serve as surrogate models for code execution prediction. To systematically investigate it, we introduce SURGE, a comprehensive benchmark with problems covering key aspects: multi-language programming tasks, competition-level programming problems, repository-level code analysis, high-cost scientific computing, time-complexity-intensive algorithms, buggy code analysis, programs dependent on specific compilers or execution environments, and formal mathematical proof verification. Through extensive analysis of open-source and proprietary LLMs, we examine scaling laws, data efficiency, and predictive accuracy. Our findings reveal important insights about the feasibility of LLMs as efficient surrogates for computational processes. The benchmark and evaluation framework are available at https://github.com/Imbernoulli/SURGE.
SURGE: On the Potential of Large Language Models as General-Purpose Surrogate Code Executors
Bohan Lyu1 * † Siqiao Huang2 * Zichen Liang1 * Qi-An Sun1 Jiaming Zhang1 1Department of Computer Science and Technology, Tsinghua. 2Institute for Interdisciplinary Information Sciences (IIIS), Tsinghua. lyubh22@gmail.com, {huang-sq23, liang-zc22}@mails.tsinghua.edu.cn *Equal contribution †Corresponding author
@@ 1 Introduction
Neural surrogate models (Zhang et al., 2024; Sun and Wang, 2019) are powerful tools in data mining and machine learning, which efficiently approximate complex computational processes. Meanwhile, Large language models (LLMs) (Reid et al., 2024; Meta, 2024; Anthropic, 2024b; Hui et al., 2024; Bi et al., 2024) have demonstrated remarkable capabilities in code-related tasks (Lu et al., 2021a; Zheng et al., 2023; Luo et al., 2023; Team, 2024a; Guo et al., 2024), including code understanding (Ahmad et al., 2020; Chakraborty et al., 2020) and code generation (Li et al., 2018a; Parvez et al., 2018). However, an equally important yet underexplored question is whether LLMs can serve as general-purpose surrogate code executors, which predict the behavior of a program without actually running it. A recent study (Lyu et al., 2024) acknowledges its importance, however, it focuses on a case study rather than a systematic analysis.
@@
The ability to predict code execution outcomes without execution has tremendous significance. In scientific computing, running simulations often requires substantial computational resources and is time-consuming, making it impractical to test every possible configuration (Lu and Ricciuto, 2019; Hesthaven and Ubbiali, 2018; Benner et al., 2015). In security-sensitive environments, executing untrusted code poses inherent risks, necessitating alternative mechanisms for assessing program behavior without exposing the system to potential vulnerabilities (Nebbione and Calzarossa, 2023; Shirazi et al., 2017; Wang et al., 2024). Additionally, some code requires highly specific execution environments, which may not always be available, making surrogate execution a valuable alternative (Queiroz et al., 2023; Gu et al., 2025). Moreover, accurately predicting a model’s potential outputs or errors is crucial for improving traditional tasks such as code understanding, code generation, and even math reasoning (Li et al., 2025). Lastly, many works use LLMs as reward models (RMs) in reinforcement learning. For code tasks, accurate execution prediction is key to a reliable RM (Ouyang et al., 2022).
@@
Traditional approaches to surrogate code executing (King, 1976; Cadar and Sen, 2013) struggle to generalize across languages and suffer from scalability issues when applied to complex real-world codebases. Containerized environments (Merkel, 2014) mitigate dependency issues but still require full code execution. Recent efforts to train neural executors (Yan et al., 2020) focus on narrow tasks and lack the generality needed for real-world code. In contrast, LLMs’ capacity to internalize patterns from vast code corpora (Lu et al., 2021b; Chaudhary, 2023) suggests a path toward general-purpose surrogate code execution.
@@
To understand the potential of LLMs as GEneral-purpose SURrogate code executors, we introduce SURGE. It includes components: (1) fundamental programming tasks in multiple languages, (2) competition programming problems requiring deep logical inference, (3) repository-level codebases that test long-range dependencies, (4) scientific simulations and optimizations where direct execution is high-cost, (5) time-consuming logical algorithms that have high time-complexity, (6) buggy code that examines LLMs’ ability to predict runtime errors, (7) programs whose behavior depends on specific compiler versions or execution environments and (8) math theorem proving in formal language (De Moura et al., 2015; Moura and Ullrich, 2021) which expects compilers to testify.
@@
Through extensive evaluation of 21 open-source and proprietary LLMs on SURGE, we provide the first large-scale study of LLMs’ capabilities as computational surrogates. Additionally, we investigate the impact of various factors, including prompt engineering strategies, programming language characteristics, computational complexity, and execution time requirements, on surrogate performance. Our findings reveal both the promising potential and current limitations of LLMs as general-purpose code execution surrogates. The performance of typical models on SURGE is shown in Figure 1.
@@
Beyond benchmarking, we conduct a scaling law study on whether LLMs’ performance improves with model size and training data. We train models with 4 different sizes on different scales of training data from the formal language subset of SURGE. Our experiments demonstrate that models’ performance consistently improves with both model size and training steps, with larger models showing stronger learning capacity and higher performance ceilings throughout the training process (Figure 2).
@@
We introduce SURGE, the first holistic benchmark for evaluating LLMs as general-purpose surrogate code executors. It consists of subsets and problems.
We evaluate open-source and proprietary LLMs on SURGE and conduct the first large-scale analysis on them.
We present a scaling law study with models of varying sizes and scales of training data, providing empirical insights on the scaling law of LLMs on these tasks.
@@ 2 Related Works
Neural surrogate models are neural network-based approximations used to replace computationally expensive simulations in various scientific and engineering domains (Zhang et al., 2024; Sun and Wang, 2019). These models act as domain-specific emulators by learning complex input-output relationships from high-fidelity data(Raissi et al., 2020; Sun et al., 2020; Bessa et al., 2017; Thuerey et al., 2020; Raissi et al., 2019; Willard et al., 2022). Recently, generative models have been incorporated into surrogate modeling. Some equip language models with traditional surrogate models to facilitate iterative optimization (Ma et al., 2024; Lyu et al., 2025), and some use generative models to realize the end-to-end surrogate process (Gruver et al., 2024; Hao et al., 2024; Wimmer and Rekabsaz, 2023; Che et al., 2024). While these studies primarily focus on natural sciences, time series, and multimodal gaming, the application of surrogate modeling to code execution, where both input and output exist in the modality of language, remains unexplored.
@@
LLMs are widely used in code-related tasks (Lu et al., 2021a; Luo et al., 2023; Team, 2024a; Guo et al., 2024), which can be fundamentally categorized into code understanding and code generation. Code understanding tasks include code summarization Hu et al. (2018); Harer et al. (2019); Ahmad et al. (2020), bug detection Li et al. (2018b); Russell et al. (2018); Zhou et al. (2019); Chakraborty et al. (2020), duplication detection Zhang et al. (2019); Yu et al. (2019); Wang et al. (2020), code retrieval (Husain et al., 2020; Lu et al., 2021a), etc. Code generation tasks include code completion Li et al. (2018a); Parvez et al. (2018), code repair Chen et al. (2019); Chakraborty et al. (2020); Lutellier et al. (2020), test generation (Watson et al., 2020; Siddiq et al., 2024; Schäfer et al., 2023), etc. However, while the potential execution result of code is important for both code understanding and generation, this aspect remains largely unexplored (Weber et al., 2024).
@@ 3 SURGE
SURGE assesses the model’s ability to approximate execution results across multiple dimensions, including multi-lingual diversity, repository-level complexity, computational intensity, error handling, and scenario-dependent variability. Below, we describe each component of SURGE, including motivation and construction methods.
| Subset | Construction Method | Source | Quantity | Metric | Categories |
| ML | Iterative Refactor | McEval (Chai et al., 2024) | 150 | Exact Match | Languages |
| CL | Iterative Refactor | GitHub | 150 | Exact Match | Difficulties & Languages |
| RL | Repository Sampling | GitHub & Custom | 60 | Mixed | - |
| SC | Manual Implementation | Custom | 150 | Mixed | Scenarios |
| TC | Manual Implementation | Custom | 150 | Mixed | Scenarios, CPU Time |
| BG | Iterative Refactor | DebugBench (Tian et al., 2024) | 150 | Jaccard similarity | Error Type & Languages |
| DR | Manual Implementation | Custom | 200 | Jaccard similarity | Variable Type |
| FL | Inference & Verification | Lean-Workbook (Ying et al., 2024) | 150 | Custom | - |
@@ 3.1 Dataset Construction
As illustrated in Figure 3, the construction of SURGE involves four distinct methodologies, each applied to specific components of our eight subsets. For ML, CL, BG components, we employ the Iterative Refactor methodology, where we refine the code through an interactive process involving LLM assistance and human verification. The RL component is constructed through Repository Sampling, where we extract and construct test cases from both public and custom code repositories. For SC, TC, DR components, we utilize Manual Implementation, carefully handcrafting code based on selected textbook materials and question collections. Finally, the FL component is developed using the Inference & Verification approach, leveraging formal mathematical provers to generate proofs and validate them through compiler verification.
To mitigate the risk of data contamination and answer leakage, we implemented a robust two-fold sanitization strategy across the dataset: (1) We applied automated filtering scripts to systematically remove comments from all code snippets before presenting them to the models for evaluation. (2)We manually inspected all generated or collected code to identify and remove any comments, ‘assert‘ statements, or other annotations that could inadvertently provide hints or reveal the expected output.
@@ 3.2 Dataset Components
A fundamental nature of a general-purpose surrogate executor is its ability to handle multiple programming languages, especially computational languages. Our dataset covers 7 such languages, including C, C++, C#, Java, Rust, Python, and Julia. Our dataset is adapted from McEval (Chai et al., 2024). The original dataset does not provide executable code, so we used an LLM to generate executable code by providing it with prompts, ground truth, and test cases in the original dataset. This generated code was then manually processed; specifically, we (1) modified code that failed to compile (e.g., by adding missing headers) and (2) carefully reviewed the code to prevent answer leakage through assert statements or comments.
@@
Next, we consider competition-level code, which presents a higher level of coding difficulty. We collect these tasks from 2 public repositories 111https://github.com/azl397985856/leetcode222https://gitee.com/shanire/OJCode, which contain problems from open coding platforms (e.g. LeetCode, Luogu). The dataset includes problems in 3 languages, C++, Java, and JavaScript. Since the original repositories only provide partial solutions, we first use an LLM to generate complete, executable code that prints the expected output. This generated code is then manually verified. During this process, if issues such as package mismatches or syntax errors (e.g., mixing language versions) arose, we manually revised the code to ensure successful compilation and execution. To investigate whether problem difficulty affects the performance of surrogate models, we classify problems into different difficulty levels.
@@
Scientific computing has long been adopting neural surrogate models. We introduced tasks ranging from solving ordinary differential equations (ODEs) to optimization problems and signal processing. These tasks are motivated by and widely used in real-world scientific challenges, including areas where increasing research has been done on solving these computational tasks through building efficient surrogate models (Wu et al., 2023; Zhang et al., 2023). A comprehensive overview of the setup for each task, along with the corresponding algorithms, can be found in Appendix C.4.1.
@@
Surrogate models were originally motivated by real-world applications where program execution is high-cost and time-consuming. It’s a necessity for LLMs to generalize well to strongly computation-power-dependent and time-consuming tasks. We include examples from linear algebra, sorting, searching, Monte Carlo simulations, and string matching programs, ensuring a broad representation of computationally intense tasks. These tasks cover various complexity classes, including P (e.g., sorting an array), NP (e.g., Hamiltonian Cycle), and NP-Hard (e.g., Traveling Salesman’s Problem). Additionally, we record the CPU execution time for each program under consistent environments individually to support subsequent studies. A detailed description of this subset is provided in Appendix C.5.1.
@@
Real-world code execution often encounters errors, which pose risks in sensitive scenarios. Therefore, code surrogate models should recognize the presence of bugs. This dataset is adapted from DebugBench (Tian et al., 2024), which extracted Java, Python, and C++ code from LeetCode and manually inserted errors from 4 major bug categories and 18 minor types. Since DebugBench only provides code snippets rather than complete executable programs, we first used an LLM to automatically complete the error-free code into fully runnable versions. After verifying their correctness, we replaced relevant parts with buggy code and executed them again to capture the corresponding error outputs. Some errors resulted in infinite loops, causing timeouts, so we set a -second execution filter to avoid such cases.
@@
Various contextual factors, such as compiler versions, optimization levels, and language standards, often influence code execution. These variations can lead to different outputs for the same code snippet. We focus specifically on C++ and manually collect code snippets from textbooks and online sources (Bryant and O’Hallaron, 2010; Lippman et al., 2012) that exhibit different behaviors under varying compilation settings. We consider multiple compilers (g++, clang++), C++ standards (03, 11, 14, 17), and optimization levels (-O0, -O1, -O2, -O3, -Os). Each snippet is executed across these different settings, and we retain only those that produce varying outputs through different configurations while discarding cases that yield identical results across all settings.
@@
Math-Proving Formal Languages are specialized programming languages designed for mathematical proof verification through compilers (De Moura et al., 2015; Moura and Ullrich, 2021; Paulson, 1994; Barras et al., 1997). These compilers can determine whether a proof is correct and identify specific errors. The verification is very time-consuming. In this study, we focus on Lean4 which is the most widely used proof language. To build our dataset, we use Goedel-Prover (Lin et al., 2025) to conduct large-scale reasoning on Lean-Workbook (Ying et al., 2024) and extract an equal proportion of correct and incorrect proofs. This balanced dataset allows us to evaluate the surrogate model’s ability to assess proof validity effectively.